Four times the smallest of three consecutive odd integers is five more than three times the
largest. What is the sum of these three consecutive odd integers? Iv'e tried really hard to get this, but even when I use consecutive numbers 1,3,5, I still can't figure it out with out getting one of the correct answers.

Found 2 solutions by ilana, stanbon:
Answer by ilana(307)

Use algebra instead of guessing and checking. Say the smallest is x. Then the 2 next odd numbers are x+2 and x+4. So you can write an equation:
4x = 3(x+4) + 5
4x = 3x + 12 + 5
4x = 3x + 17
x = 17
So the integers are 17, 19, and 21.
The sum of these integers is 57.

Answer by stanbon(75887)

Four times the smallest of three consecutive odd integers is five more than three times the largest.
What is the sum of these three consecutive odd integers?
---------------------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
---------
Equation:
4(2x-1) = 3(2x+3)+5
-----
8x-4 = 6x+14
2x = 18
x = 9
-----
1st: 2x-1 = 17
2nd: 2x+1 = 19
3rd: 2x+3 = 21
----------------------
The sum of these numbers = 6x+3 = 57
===============================
Cheers,
Stan H.
===========